OpenStudy (ranga):

a^3 - 3a^2 + 5a = 1 b^3 - 3b^2 + 5b = 5 a + b = ?

OpenStudy (freckles):

\[a^3-b^3-3a^2+3b^2+5a-5b=1-5 \\ (a-b)(a^2+ab+b^2)-3(a-b)(a+b)+5(a-b)=-4 \\ (a-b)[(a^2+ab+b^2)-3(a+b)+5]=-4 \\ (a-b)(a^2+ab+b^2-3a-3b+5)=-4 \\ \] So far this is what I thought up.. I'm not sure if I will continue on this path or change it up.

OpenStudy (ranga):

I went the other way and aded them up but did not get me anywhere.

OpenStudy (freckles):

and I guess we are looking for a+b not a-b anyways

OpenStudy (ranga):

Each has one real root and two complex roots. If you add up the corresponding roots of each equation they all add up to 2!

OpenStudy (ranga):

I know the three roots of a cubic equation, ax^3 + bx^2 + cx + d, add to -b/a. Their product is -d/a.

OpenStudy (loser66):

if it is so, the a from your last equation is 1 , b = -3, c=5 and d =-1, the original one is a^3 -3a^2 +5a -1 =0 have 2 roots a1, a2 which satisfy a1+a2= 3 and a1*a2= 1

OpenStudy (loser66):

Do I mess it up? You use the same letter to indicate to 2 different problems. hehehe... that makes me dizzy

OpenStudy (ranga):

Cubic equation have three roots. So a1 + a2 + a3 = 3; a1*a2*a3 = 1

OpenStudy (loser66):

anyway, your cubic equation don't have integer solutions. I don't know how to find them out but (a + b) = 1/ the rest of a equation + 5/ the rest of b equation I factor a , b out , isolate them, then add them up. hehehe... stupid way but I don't know how to do more.

OpenStudy (ranga):

This was an exam question where no calculators or graphing devices allowed. So they don't expect people to find the roots to answer this question.

OpenStudy (ranga):

If the two equations can be combined and expressed as follows, then we may be able to solve it: (a+b)^3 + k1(a+b)^2 + k2(a+b) + k3 = 0 then this will have one rational root, namely, a+b = 2.

OpenStudy (loser66):

too complicated. For which grade the problem is ?

OpenStudy (ranga):

This was for a college entrance exam in India that my nephew took last year.

OpenStudy (loser66):

Wow!! India is the leader on Math and computer.

OpenStudy (ranga):

This particular engineering college is tough to get into and each year they come up with questions in Math, Physics, Inorganic and Organic Chemistry that is just awesome.

OpenStudy (loser66):

Agree!! However, awesome if we can solve it. If we cannot, no more awesome, hehehe.. I think so.

OpenStudy (ranga):

Oh, this will get solved eventually. I will just forget about it and go about my business and sooner or later the answer is likely to pop up. Or someone here may be able answer this.

OpenStudy (rational):

any restrictions on a, b ? are they integers ?

Miracrown (miracrown):

Ok so let's see...have you tried solving for 'a' and 'b' simultaneously?

OpenStudy (ranga):

If a1, a2, a3 are the three roots of a^3 - 3a^2 + 5a - 1 = 0, then: a1 + a2 + a3 = 3 a1 * a2 * a3 = 1 a1*a2 + a2*a3 + a3*a1 = 5 If b1, b2, b3 are the three roots of b^3 - 3b^2 + 5b - 5 = 0, then: b1 + b2 + b3 = 3 b1 * b2 * b3 = 5 b1*b2 + b2*b3 + b3*b1 = 5

OpenStudy (ranga):

We want a1+b1 or a2 + b2 or a3 + b3.

OpenStudy (anonymous):

lol i am back at the adding step...

OpenStudy (ranga):

No restrictions given in the problem for a and b. But we cannot assume they are integers. Each equation has one real, non-integer root and two complex roots if I use a graphing tool. But it is not allowed.

OpenStudy (rational):

may be we can reverse engineer - show that the upper and lower bounds of a+b is 2

OpenStudy (freckles):

if we can some factor equ1+equ2 to into (a+b-2)(times whatever)=0 then we got it

OpenStudy (ranga):

Roots of the first equation are: 0.22908, 1.3855 + 1.5639i, 1.3855 - 1.5639i Roots of the first equation are: 1.77092, 0.6145 - 1.5639i, 0.6145 + 1.5639i Look how nicely each of the corresponding root pairs add to 2!

OpenStudy (ranga):

*second*

OpenStudy (freckles):

somehow*

OpenStudy (freckles):

but i guess that is what ranga said above somewhere anyways

OpenStudy (anonymous):

wolf implicit plot gives a line hmmm

OpenStudy (ranga):

Oh my!

OpenStudy (anonymous):

no idea why i don't thing you can write this as a polynomial in \(a+b\) because \((a+b)^3\) and terms in it like \(ab^2\) that you can't get rid of with \((a+b)^2\) or \(a+b\) so i although it looks like the right approach, i have no idea how you can carry out that plan

OpenStudy (freckles):

Like I get this so far \[(a+b)^3-3(a+b)^2+5(a+b)=3ab(a+b-2)+6\]

OpenStudy (freckles):

which is neat because the other thingys we had were in this form you know \[a^3-3a^2+5a=1 \text{ and } b^3-3b^2+5b=5\]

OpenStudy (freckles):

so we have f(a)=1 and f(b)=5 and f(a+b)=3ab(a+b-2)+6

OpenStudy (rational):

max/min of \(f : a + b \) subject to \(g : a^3+b^3-3(a^2+b^2)+5(a+b)=6\) \(\nabla f = \lambda \nabla g \) \(\langle1, 1 \rangle = \lambda \langle 3a^2 -6a + 5, 3b^2-6b+5 \rangle\) \(\implies 3a^2 - 6a + 5 = 3b^2 - 6b + 5\) \(\implies 3(a^2-b^2) - 6(a-b) = 0\) \(\implies a=b ~ \lor ~~a+b = 2\)

OpenStudy (rational):

^lagrange may be luxury for college entrance problem.. but thats all i can think of as of now..

OpenStudy (anonymous):

wish i had thought of that

OpenStudy (ranga):

\[ (1):~~~a^3 - 3a^2 + 5a = 1 \\ (2):~~~b^3 - 3b^2 + 5b = 5 \\ \text{ }\\ (1):~~~a^3 - 3a^2 + 5a = 1 \\ \text{The }a^3 - 3a^2 \text{ terms suggest } (a-1)^3.~~~~~ (a-1)^3 = a^3 - 3a^2 + 3a - 1 \\ (1):~~~(a-1)^3 + 2a + 1 = 1 \\ (1):~~~(a-1)^3 + 2(a-1) + 2 + 1 = 1 \\ (1):~~~(a-1)^3 + 2(a-1) = -2 \\\\ \text{ }\\ (2):~~~b^3 - 3b^2 + 5b = 5. ~~~~ (b-1)^3 = b^3 - 3b^2 + 3b - 1 \\ (2):~~~(b-1)^3 + 2(b-1) + 3 = 5 \\ (2):~~~(b-1)^3 + 2(b-1) = 2 \\\\ \text{ }\\ \text{Let } f(x) = x^3 + 2x \\ \text{ }\\ (1): f(a-1) = -2 \\ (2): f(b-1) = 2 \\ \text{ }\\ f(x) \text{ is an odd function. That is, } f(x) = -f(-x) \\ \text{ }\\ \text{Therefore, from (1) and (2) above, } a - 1 = -(b - 1) \\ a-1 = -b + 1 \\ a+b = 2 \\ \]

OpenStudy (rational):

beautiful !! xD

OpenStudy (freckles):

omg, ranga that is the "cutest" thing I ever seen

OpenStudy (ranga):

Cuter than kittens or puppies?! lol. Thank you all. Here is another interesting problem from the same test. This has been solved and so I will close this thread. \[\Large \text{If }~a, b, c, d ~ \in ~ I^+ \text{ and }~ \log_{a} b = \frac 32, ~ \log_{c} d = \frac 54 \text{ and } \\ \Large (a - c) = 9 \text{ then the remainder when }~b^d~\\ \text{ }\\\Large \text{is divided by 31 is _______.}\]

OpenStudy (ranga):

a,b,c,d are positive integers.

OpenStudy (freckles):

The best way I came up with his trial and error and finally by fermats little theorem I got the answer is 1.

OpenStudy (ranga):

\[ b = a^{3/2};~~~ d = c^{5/4} \\ \text{ }\\ \text{If a,b,c,d are all integers, then 'a' has to be a perfect square and 'c' has to be } \\ \text{a perfect 4th power because we have to take the square root of 'a' and the }\\ \text{fourth root of 'c' to find b and d.}\\ \text{ }\\ \text{Let }a = p^2; ~~c = q^4 \text{ where p and q are integers.}\\ a - c = p^2 - q^4 = (p - q^2)(p + q^2) = 9 = 1 \times 9 \\ p - q^2 = 1; ~~~ p + q^2 = 9 \\ 2p = 10; ~~~ p = 5; ~~~ q = 2; ~~~ a = 25; ~~~ c = 16; ~~~ b = 125; ~~~ d = 32 \\ \text{ }\\ b^d~\text{mod}~31 = 125^{32}~\text{mod}~31 = (125~\text{mod}~31)~\text{mod}~31 = 1~\text{mod}~31 = 1. \]

OpenStudy (ranga):

Yes, freckles, you got the same final answer as mine!

OpenStudy (ranga):

In the last line an exponent was left out: \[ b^d~\text{mod}~31 = 125^{32}~\text{mod}~31 = (125~\text{mod}~31)^{32}~\text{mod}~31 = 1~\text{mod}~31 = 1. \]

OpenStudy (rational):

\(a^{30}\equiv 1 \mod 31\)

OpenStudy (rational):

interestin ! did u show \(30 | \frac{3}{2} c^{\frac{5}{4}} \) @freckles ?

OpenStudy (rational):

wow ! @ranga 's solution looks very neat xD

OpenStudy (ranga):

Thank you @rational. Enjoy sharing these atypical problems.